OpenStudy (anonymous):

Can someone help me understand how to write in point-slope form an equation of the line through each pair of points (1,9) (6,2)

OpenStudy (anonymous):



OpenStudy (anonymous):

so do you know the point slope form?

OpenStudy (anonymous):

No, I really need help with the whole thing.

OpenStudy (anonymous):

recall it is \[(y-y _{1})=m(x-x _{1})\]

OpenStudy (anonymous):

does that look familiar?

OpenStudy (anonymous):

MEOW

OpenStudy (anonymous):

you first need to find m (the slope) by doing (y1-y2)/(x1-x2)

OpenStudy (anonymous):

so (9-6)/(2-1)=3/1=3

OpenStudy (anonymous):

now the\[y \] and \[x\] stays in the formula.

OpenStudy (anonymous):

so just plug in what you know, you have two points, and you only need one in your point slope formula. so you can pick either one of the point to plug in first of what you know

OpenStudy (anonymous):

wait, so do I need (1,6)=m (9,2) anywhere or is that even correct? I need a complete breakdown, my teacher's no help, sorry.

OpenStudy (anonymous):

Maggie, you don't need (1,6), you have to invert it first so it's (6,1)

OpenStudy (anonymous):

not quite, \[(y-y _{1})=m(x-x _{1})\] \[(y-9)=m(x-1)\] notice how i just got one of your points and plugged it in for \[(x _{1},y _{1})\]

OpenStudy (anonymous):

Its like this

OpenStudy (anonymous):

so now we just need to find the slope, you have two points given right? its (1,9) and (6,2)

OpenStudy (anonymous):